OpenStudy (anonymous):

The product of (a + b)(a - b) is a^2 - b^2. Answer Sometimes Always Never

OpenStudy (mertsj):

always

OpenStudy (anonymous):

Always. (a+b)(a-b) = a^2 - ab + ba - b^2 = a^2 - b^2

OpenStudy (anonymous):

It can't be ever anything else.